# The Rational Torsion Subgroup of $J_{0}(\mathfrak{p}^{r})$

### v1

@@ 6 Proof of Main Theorem
$\widetilde{r}(\Delta_{\mathfrak{p}^{k}})(\begin{pmatrix}1&0\\ \mathfrak{p}&1\end{pmatrix}\begin{pmatrix}\pi_{\infty}^{-j}&0\\ 0&1\end{pmatrix})$
$=\widetilde{r}(\Delta)(\begin{pmatrix}\mathfrak{p}^{k}&0\\ 0&1\end{pmatrix}\begin{pmatrix}1&0\\ \mathfrak{p}&1\end{pmatrix}\begin{pmatrix}\pi_{\infty}^{-j}&0\\ 0&1\end{pmatrix})=\widetilde{r}(\Delta)(\begin{pmatrix}\mathfrak{p}^{k}\pi_{\infty}^{-j}&0\\ \mathfrak{p}\pi_{\infty}^{-j}&1\end{pmatrix})$
@@
where the penultimate equality again comes from [5, Eq. (2.3)]. ∎
@@
By Prop. 6.6, $\mathcal{E}_{0}(\mathfrak{n},R)[U_{\mathfrak{p}}]$ is contained in $R\cdot E_{\mathfrak{n}}$. Suppose that $c\cdot E_{\mathfrak{n}}\in\mathcal{E}_{0}(\mathfrak{n},R)[U_{\mathfrak{p}}]$ for some $c\in R$. Let $e_{j}:=\begin{pmatrix}1&0\\ \mathfrak{p}&1\end{pmatrix}\begin{pmatrix}\pi_{\infty}^{-j}&0\\ 0&1\end{pmatrix}\in E(\mathscr{T})$ with $j\geq 0$. By Lem. 6.7, we have $c\cdot E_{\mathfrak{n}}(e_{j})=c\cdot q^{j+1}M(\mathfrak{p})$ for $j\geq 0$. Since $q\in R^{\ast}$, we have $c\cdot E_{\mathfrak{n}}(e_{j})=0$ for any $j\gg 0$ if and only if $c\in R\left[M(\mathfrak{p})\right]$, which completes the proof. ∎
@@
Let $R:=\mathbb{Z}/\ell^{k}\mathbb{Z}$ with $k\gg 0$ such that $\mathcal{T}(\mathfrak{n})_{\ell}\hookrightarrow\mathcal{E}_{00}(\mathfrak{n},R)$ in Lem. 6.2. By Lem. 6.8,
$\mathcal{E}_{00}(\mathfrak{n},R)[U_{\mathfrak{p}}]\subseteq\mathcal{E}_{0}(\mathfrak{n},R)[U_{\mathfrak{p}}]\subseteq R\left[M(\mathfrak{p})\right]\cdot E_{\mathfrak{n}}\cong\frac{\mathbb{Z}_{\ell}}{M(\mathfrak{p})\mathbb{Z}_{\ell}}.$
@@
From the above, we obtain Main Theorem.

### v2

@@ 6 Proof of the Main Theorem
$\widetilde{r}(\Delta_{\mathfrak{p}^{k}})(e_{j})=\widetilde{r}(\Delta)(\begin{pmatrix}\mathfrak{p}^{k}&0\\ 0&1\end{pmatrix}e_{j})=\widetilde{r}(\Delta)(\begin{pmatrix}\mathfrak{p}^{k}\pi_{\infty}^{-j}&0\\ \mathfrak{p}\pi_{\infty}^{-j}&1\end{pmatrix})$
@@
where the penultimate equality again comes from [5, Equation (2.3)]. ∎
@@
By Proposition 6.6, $\mathcal{E}_{0}(\mathfrak{n},R)[U_{\mathfrak{p}}]$ is contained in $R\cdot E_{\mathfrak{n}}$. Suppose that $c\cdot E_{\mathfrak{n}}\in\mathcal{E}_{0}(\mathfrak{n},R)[U_{\mathfrak{p}}]$ for some $c\in R$. Let $e_{j}:=\begin{pmatrix}1&0\\ \mathfrak{p}&1\end{pmatrix}\begin{pmatrix}\pi_{\infty}^{-j}&0\\ 0&1\end{pmatrix}\in E(\mathscr{T})$ with $j\geq 0$, then $E_{\mathfrak{n}}(e_{j})=q^{j+1}M(\mathfrak{p})$ by Lemma 6.7. Since $q\in R^{\ast}$, we have $c\cdot E_{\mathfrak{n}}(e_{j})=0$ for any $j\gg 0$ if and only if $c\in R\left[M(\mathfrak{p})\right]$, which completes the proof. ∎
@@
Let $R:=\mathbb{Z}/\ell^{k}\mathbb{Z}$ with $k\gg 0$ such that $\mathcal{T}(\mathfrak{n})_{\ell}\hookrightarrow\mathcal{E}_{00}(\mathfrak{n},R)$ in Lemma 6.2. By Lemma 6.8,
$\mathcal{E}_{00}(\mathfrak{n},R)[U_{\mathfrak{p}}]\subseteq\mathcal{E}_{0}(\mathfrak{n},R)[U_{\mathfrak{p}}]\subseteq R\left[M(\mathfrak{p})\right]\cdot E_{\mathfrak{n}}\hookrightarrow\frac{\mathbb{Z}_{\ell}}{M(\mathfrak{p})\mathbb{Z}_{\ell}}.$
@@
From the above, we obtain our Main Theorem.
7 Acknowledgments
The author sincerely dedicates this paper to his supervisor, Prof. Mihran Papikian, for his invaluable guidance and encouragement. Special thanks are also extended to Prof. Fu-Tsun Wei for the insightful discussions on modular units. The author also expresses gratitude to Prof. Chieh-Yu Chang for the invitation to present this paper at NTHU and to Prof. Wen-Ching Winnie Li for the invitation to collaborate on a survey paper on Ogg’s conjectures. Lastly, heartfelt appreciation goes to Chiu-Lin Huang for her companionship and emotional support.